OpenStudy (anonymous):

\[\sqrt{90} + \sqrt{250}\]

OpenStudy (skullpatrol):

Any ideas?

OpenStudy (skullpatrol):

9 = 3*3 25 = 5*5

OpenStudy (anonymous):

I get that 9 x 10 = 90 and 25 x 10 = 250 but how the hell did they get 8 and 10

OpenStudy (anonymous):

so i break the 9 and 25 further and combine them?

OpenStudy (skullpatrol):

yup

OpenStudy (anonymous):

could you draw it out so i know im doing the correct steps

OpenStudy (skullpatrol):